# Kekulé order from diffuse nesting near higher-order Van Hove points

### v1

@@ Abstract
Translation symmetry-breaking order is assumed to be suppressed by the lack of Fermi surface nesting near certain higher-order Van Hove singularities (HOVHS). We show the anisotropic band-flattening inherent to such HOVHS, combined with broadening of the Fermi surface due to elevated critical temperatures, results in the Fermi surface becoming approximately nested at a wavevector unrelated to the precise shape of the Fermi surface—leading to a $\sqrt{3}\times\sqrt{3}$ Kekulé density wave formation. The effect is demonstrated using unbiased renormalization group calculations for a model of the breathing kagome lattice. Our mechanism—termed diffuse nesting—represents an entirely new notion in the study of Fermi surface instabilities.
@@
From a technical perspective, the transition from a logarithmic to an algebraic divergence of the DOS substantially alters the weak coupling picture for HOVHS, as $\chi_{pp}$ and $\chi_{ph}$ likewise attain algebraic divergences. For many such saddle points, the anisotropy of the band flattening suppresses the nesting condition, rendering the Fermi liquid unstable to translationally invariant particle-hole instabilities at zero momentum transfer, rather than density wave order, i.e., $\bm{q}=0$ Pomeranchuk instabilities such as nematic order or altermagnetism [86, 87, 71, 88].
@@
In this Letter, we propose an unprecedented mechanism for the formation of translation-symmetry breaking $\sqrt{3}\times\sqrt{3}$ density waves in hexagonal lattice systems with HOVHS. The phenomenon arises from the interplay between warped Fermi surfaces induced by anisotropic HOVHS and elevated critical scales, leading to what we term diffuse nesting. That is, the associated nesting condition manifests only when the Fermi surface retains a finite broadening, e.g. due to finite temperature, or potentially disorder. From an itinerant perspective, a high density of states implies elevated critical temperatures of ordered states, which makes this condition natural in systems with HOVHS. Remarkably, the resulting instabilities are characterized by a wavevector $\bm{q}=K$ unrelated to the momenta connecting the HOVHS at the $M$-point, and so do not involve the regions of the Fermi surface where the density of states is largest.
@@
The proposed state is reminiscent of Kekulé order, first conceived as the ground state of Benzene. It describes an alternating hexagonal $\sqrt{3}\times\sqrt{3}$ bond order, previously discussed in the context of carbon nanotubes, graphene [89, 90, 91, 92, 93, 94, 95], and more recently twisted graphene multilayers [96, 97, 98, 99, 100, 101, 102]. The similarity to the Kekulé orders previously studied in honeycomb systems is emphasized by examining the line graph reconstruction of the kagome lattice, which reveals a Kekulé-Y pattern on the resulting effective honeycomb structure.
@@
Our findings are illustrated by functional renormalization group calculations [103, 104, 105, 106, 107, 108, 109] for a model inspired by the breathing kagome surface of ferromagnetic Co3Sn2S2, which features HOVHS near the Fermi level [71]. We trace the origin of the $K$-ordering tendency to specific properties inherent to such HOVHS and propose a simplified single-band model, featuring an eigenspectrum solely determined by series-expanded saddle points situated at the $M$-points—thus generalizing our observation to arbitrary hexagonal lattice systems. While we focus on a model of spinless fermions, our results can be extended to spinful systems, revealing magnetic Kekulé-type order in the breathing kagome configuration, as detailed in the Supplemental Material (SM) [110].
@@
Many prior studies of higher-order Van Hove points have worked with so-called $g$-ology or patch models, which approximate the Fermi surface as isolated points at the VHS [86, 87, 111, 112, 113]. Our results dramatically illustrate that such models necessarily fail to capture certain ordering tendencies of a HOVHS, which counter-intuitively involve the fermions away from the momenta where the density of states is largest. The phenomenon we describe represents an entirely new notion in the study of Fermi surface instabilities.
Model.
Figure 1: Breathing kagome model. (a) Breathing kagome lattice configuration with hoppings up to fourth nearest neighbor. (b) Corresponding band structure with density of states, characterized by an algebraic divergence at the Fermi level. The inset displays the broadened Fermi surface at $T=0.01\,t$.
@@
HOVHS can emerge in tight-binding models when longer-ranged or complex hoppings are introduced. In parameter regimes where the quadratic term of the dispersion near the Van Hove point is suppressed, a quartic dispersion emerges [86, 87]. One concrete realization is the breathing configuration of the kagome Co3Sn surface termination (Fig. 1a) of ferromagnetic shandite compound Co3Sn2S2 [71]. The tight-binding model adapted from ab initio calculations reads
@@
where $c_{i}$ annihilates an electron at site $i$, $n_{i}=c_{i}^{\dagger}\,c^{\phantom{\dagger}}_{i}$ is the electron density operator and $\mu$ the chemical potential; $t_{1\rightarrow 4}\approx(-1.383,-0.617,0.352,-0.176)\,t$ correspond to the hopping amplitudes up to fourth nearest neighbors with the energy scale $t$, see Fig. 1a. The hopping $t_{4}$ vanishes along a particular direction for each sublattice, indicated as $\tilde{t}_{4}$ in Fig. 1a. In addition to the kinetic term, we add repulsive density-density interactions $V_{n}$.
The resulting dispersion, which features a HOVHS at the $M$-points, is shown in Fig. 1b, along with the density of states (DOS) hosting an algebraic divergence at the Fermi level; a notable difference can be observed between the DOS of the HOVHS and that of the ordinary VHS near $\varepsilon\approx 6t$ in the upper band. An additional HOVHS emerges at $\bm{q}=\Gamma$, where the two upper bands touch. The inset displays the Fermi surface (FS) for a finite window of $\mu$. Due to the anisotropic higher-order nature of the dispersion along the $\Gamma M$ direction—which becomes proportional to $\alpha\,k^{2}_{KM}-\beta\,k^{4}_{\Gamma M}$, where $k_{ab}$ denotes the momentum component along direction $ab$—the FS becomes significantly flattened. As a consequence, the perfect nesting at transfer momentum $M_{i}$, which is typical for ordinary VHS in hexagonal lattice systems, is lifted. Although the majority of the DOS remains concentrated at the $M$-points, the band flattening induces an enhanced spread along the Fermi surface (FS) surrounding $M$.
@@
Figure 2: Leading eigenvalues and order parameter. (a) Leading eigenvalues of the effective scattering vertex in the particle-hole channel on a transfer momentum $\boldsymbol{q}$-mesh in the Brillouin zone. The dominant eigenvalues are situated at the two inequivalent $K$-points. (b) Order parameter obtained by combining functional renormalization group calculations with subsequent Landau-Ginzburg and mean-field analysis.
While the pristine kagome Hubbard model, featuring perfect nesting and sublattice interference at its Van Hove fillings, has been studied extensively using FRG methods [6, 115, 114], the influence of HOVHS has mainly been investigated either by mean-field theories [78], or by renormalization group approaches applied to simple analytical models, using Fermi surface patching schemes [86, 87, 113, 112]. Relying on the assumption that the relevant fermions are those where the density of states is largest, these simple models discard the Fermi surface away from the vicinity of the HOVHS. For HOVHS near $M$, the FS scattering processes are then restricted to transfer momenta $\boldsymbol{q}=(M,\,\Gamma)$. However, by relaxing this constraint within our FRG analysis, we observe an unprecedented ordering tendency towards $\bm{q}=K$ density waves—signified by the dominant peaks at $K$ in Fig. 2a—with significant bond character, i.e., Kekulé order. Even though this wavevector is unrelated to the position of the HOVHS, we observe the Kekulé instability in an extensive parameter regime (see SM [110]).
This instability of the particle-hole channel breaks translation symmetry, and corresponds to the two inequivalent momenta $K$ and $K^{\prime}$ related by time-reversal symmetry (Fig. 2a). The enlarged $\sqrt{3}\times\sqrt{3}$ periodicity triples the size of the unit cell (Fig. 2b, black). The resulting order parameter is given by the eigenstates of the two-particle vertex, and is found to correspond to the $E^{\prime}_{3}$ irreducible representation of the extended point group $C^{\prime\prime}_{3v}$ (see SM [110]).
@@
$\Delta_{\boldsymbol{k}}(A,\phi)=A\,\big{[}\mathrm{e}^{\mathrm{i}\phi}\Delta_{\boldsymbol{k}}^{K}+\mathrm{e}^{-\mathrm{i}\phi}\Delta_{\boldsymbol{k}}^{{K^{\prime}}}\big{]}\,,$ (2)
@@
It is noteworthy that our kagome Kekulé phase exhibits real-space features similar to those previously reported in honeycomb models. The VHS near the Fermi level in Co3Sn2S2 is the remnant of the so-called mixed ($m$)-type VHS in the undistorted phase [71], near which the wavefunction at each $M$-point resides on two of the three kagome sublattices. The local density of states for such a Fermi surface forms a honeycomb lattice situated in the center of the up and down triangles [117]. Upon triggering a breathing distortion, the density of states depletes on one sublattice of this emergent honeycomb lattice; the result is that the Fermi level density of states near the HOVHS forms a partially sublattice-imbalanced honeycomb lattice, as can be seen in the nematic density of states simulated in Ref. [71]. In real-space, our Kekulé state is therefore expected to closely resemble the Kekulé orders found in honeycomb systems, but with an imbalance between the $A$ and $B$ sublattices. This correspondence can also be visualized through the line graph reconstruction of the kagome lattice, in which the lattice sites can be associated with the bonds of the emergent honeycomb lattice. This construction reveals a Kekulé-Y order and implies that the state given in Fig. 2b transforms in the same irreducible representation (see SM [110]).
Diffuse nesting.
Figure 3: Susceptibility and Fermi surface nesting. (a) Maximal eigenvalue of the particle-hole susceptibility Eq. 3 for three different temperatures $T$. At higher $T$, the peak at momentum $K$ surpasses the $\Gamma$ peak. (b) Overlap (black) of initial (gray) and $K$-shifted (purple) Fermi surface for $T=10^{-2}\,t$. Circular nesting features drive $K$-order at finite $T$.
The emergence of $K$-order across a broad interaction parameter regime indicates that the initial structure of the interaction vertex is largely irrelevant to the underlying ordering mechanism, implying that its origin can be discerned from a kinetic viewpoint, i.e., the identification of a generalized nesting condition. We consider the static particle-hole susceptibility at finite momentum and temperature,
@@
where $G_{o_{1}o_{2}}(\bm{k},\omega_{n})$ is the single particle propagator with momentum $\bm{k}$ and fermionic Matsubara frequency $\omega_{n}$ and $o_{i}$ indicates the sublattice degree of freedom. Figure 3a shows that at low temperature, the susceptibility peak at $\bm{q}=\Gamma$ dominates—in line with the expectation that translationally invariant orders should dominate at weak coupling—yet at a finite temperature $T$, the susceptibility at the transfer momentum $\bm{q}=K$ exceeds that at $\bm{q}=\Gamma$.
@@
One can understand Fermi surface nesting as the overlap of the FS with itself upon shifting by a momentum $\bm{q}$. For an unbroadened FS, i.e. in the limit of $T\rightarrow 0$, no significant overlap can be discerned for nonzero $\boldsymbol{q}$ nesting configurations, owing to the different curvatures of overlapping portions of the FS. However, finite temperature $T$ can be understood as a broadening of the FS, as implied by Fermi-Dirac statistics smearing out the states near the Fermi level (Fig. 1b, inset). For an exemplary value of $T=10^{-2}\,t$, Fig. 3b shows nested circular regions (black) that emerge when the broadened FS is shifted by $\boldsymbol{q}=K$, revealing the origin of the dominant finite-momentum peak in the susceptibility displayed in Fig. 3a. A detailed discussion and visualization for various temperatures and nesting vectors is provided in the SM [110].
@@
We emphasize that this argument likely generalizes beyond our finite temperature calculations, and is expected in the presence of other kinds of Fermi surface broadening, such as that due to interactions or disorder. Finite interactions of intermediately correlated systems elevate the critical temperature and allow FS scattering processes to access an increasing energy window around the Fermi level comparable to the broadening in Fig. 3b. As a consequence, we observe $K$-order across the predominant part of the finite interaction parameter regime within our FRG analysis (see SM [110]).
@@
Previous studies suggest enhanced nematic ordering tendencies associated with HOVHS [86, 87, 71]. These prior models are valid in the strict weak coupling regime, and indeed we observe competing $\boldsymbol{q}=\Gamma$ Pomeranchuck fluctuations—reflected by the peak at $\Gamma$ in Fig. 2a—which become dominant for very weak interaction parameters. Note that the enhanced DOS of HOVHS effectively places the system in an intermediate-coupling regime even for relatively small interactions—further disfavoring the applicability of previous approaches. This renders the functional renormalization group—valid from the weakly to intermediately coupled limits—a vital tool to appropriately capture the leading ordering tendency.
@@
correspondingly decreases, allowing for more localized Wannier functions [119, 120] (see SM [110] for a quantitative illustration). This allows for a conceptually generalized model consisting of a piecewise series-expanded dispersion around the $M$-points of a hexagonal BZ with quartic dispersion along the $\Gamma M$ direction, as visualized in Fig. 4a. The shaded triangular regions represent the sections of the BZ where the individual expansions govern the dispersion. This piecewise definition preserves the lattice symmetry and ensures a continuous spectrum [110]. The resulting bandstructure of the effective single orbital model and its FS are displayed in Fig. 4b. Apart from the cusps at the $\Gamma K$ line, the FS retains the same directional band flattening features around the $M$-points. Consequently, diffuse nesting persists, and the dominance of $K$-order is observed within our unbiased FRG analysis, generalizing our proposed diffuse nesting mechanism to arbitrary lattice systems featuring such HOVHS.
@@
SM 1 Kinetic models with higher-order Van-Hove singularities

### v2

@@ Abstract
Translation symmetry-breaking order is assumed to be suppressed by the lack of Fermi surface nesting near certain higher-order Van Hove singularities (HOVHS). We show the anisotropic band-flattening inherent to such HOVHS, combined with broadening of the Fermi surface due to elevated critical scales, results in the Fermi surface becoming approximately nested at a wavevector unrelated to the precise shape of the Fermi surface—leading to a $\sqrt{3}\times\sqrt{3}$ Kekulé density wave formation. The effect is demonstrated using unbiased renormalization group calculations for a model of the breathing kagome lattice. Our mechanism—termed diffuse nesting—represents an entirely new notion in the study of Fermi surface instabilities.
@@
From a technical perspective, the transition from a logarithmic to an algebraic divergence of the DOS substantially alters the weak coupling picture for HOVHS, as $\chi_{pp}$ and $\chi_{ph}$ likewise attain algebraic divergences. For many such saddle points, the anisotropy of the band flattening suppresses the nesting condition, rendering the Fermi liquid unstable to translationally invariant particle-hole instabilities at zero momentum transfer, rather than density wave order, i.e. $\bm{q}=0$ Pomeranchuk instabilities such as nematic order or altermagnetism [21, 34, 74, 85].
@@
In this Letter, we propose an unprecedented mechanism for the formation of translation-symmetry breaking $\sqrt{3}\times\sqrt{3}$ density waves in hexagonal lattice systems with HOVHS. The phenomenon arises from the interplay between warped Fermi surfaces induced by anisotropic HOVHS and elevated critical scales, leading to what we term diffuse nesting. That is, the associated nesting condition manifests only when the Fermi surface retains a finite broadening, e.g. due to finite temperature, interactions, or potentially disorder. From an itinerant perspective, a high density of states implies elevated critical scales of ordered states, which makes this condition natural in systems with HOVHS. Remarkably, the resulting instabilities are characterized by a wavevector $\bm{q}=K$ unrelated to the momenta connecting the HOVHS at the $M$-point, and so do not involve the regions of the Fermi surface where the density of states is largest.
@@
The resulting state is a variant of Kekulé order, first conceived as the ground state of Benzene. It describes an alternating hexagonal $\sqrt{3}\times\sqrt{3}$ bond order, previously discussed in the context of carbon nanotubes, graphene [73, 12, 38, 77, 16, 120, 30], and more recently twisted graphene multilayers [108, 60, 112, 116, 45, 61, 119]. The similarity to the Kekulé orders previously studied in honeycomb systems is emphasized by examining the line graph reconstruction of the kagome lattice, which reveals a Kekulé-Y pattern on the resulting effective honeycomb structure.
@@
Our findings are illustrated by functional renormalization group calculations [97, 71, 87, 90, 5, 89, 88] for a model inspired by the breathing kagome surface of ferromagnetic Co3Sn2S2, which features HOVHS near the Fermi level [74]. We trace the origin of the $K$-ordering tendency to specific properties inherent to such HOVHS and propose a simplified single-band model, featuring an eigenspectrum solely determined by series-expanded saddle points situated at the $M$-points—thus generalizing our observation to arbitrary hexagonal lattice systems. While we focus on a model of spinless fermions—in which any density wave necessarily occurs in the charge channel—our results can be extended to spinful systems, revealing magnetic Kekulé-type order in the breathing kagome configuration, as detailed in the Supplemental Material (SM) [106].
@@
Many prior studies of higher-order Van Hove points have worked with so-called $g$-ology or patch models, which approximate the Fermi surface as isolated points at the VHS [21, 34, 11, 123, 39]. Our results dramatically illustrate that such models necessarily fail to capture certain ordering tendencies of a HOVHS, which counter-intuitively involve the fermions away from the momenta where the density of states is largest, reminiscent of the imperfect nesting scenario in Ref. [52]. The phenomenon we describe represents an entirely new notion in the study of Fermi surface instabilities.
@@
Model.—HOVHS can emerge in tight-binding models when longer-ranged or complex hoppings are introduced. In parameter regimes where the quadratic term of the dispersion near the Van Hove point is suppressed, a quartic dispersion emerges [21, 34]. One concrete realization is the breathing configuration of the kagome Co3Sn surface termination (Fig. 1a) of ferromagnetic shandite compound Co3Sn2S2 [74]. The tight-binding model adapted from ab initio calculations reads
@@
where $c_{i}$ annihilates an electron at site $i$, $n_{i}=c_{i}^{\dagger}\,c^{\phantom{\dagger}}_{i}$ is the electron density operator and $\mu$ the chemical potential; $t_{1\rightarrow 4}\approx(-1.383,-0.617,0.352,-0.176)\,t$ correspond to a choice of hopping amplitudes up to fourth nearest neighbor with energy scale $t$, producing an exact HOVHS (Fig. 1a). These parameters deviate slightly from the fitted ab initio calculation, see Section A and Section B [106] for details and stability analyses. The hopping $t_{4}$ vanishes along a particular direction for each sublattice, indicated as $\tilde{t}_{4}$ in Fig. 1a. In addition to the kinetic term, we add repulsive density-density interactions $V_{n}$.
The resulting dispersion, which features a HOVHS at the $M$-points, is shown in Fig. 1b, along with the density of states (DOS) hosting an algebraic divergence at the Fermi level; a notable difference can be observed between the DOS of the HOVHS and that of the ordinary VHS near $\varepsilon\approx 6t$ in the upper band. An additional HOVHS emerges at $\bm{q}=\Gamma$, where the two upper bands touch. The inset displays the Fermi surface (FS) for a finite window of $\mu$. Due to the anisotropic higher-order nature of the dispersion along the $\Gamma M$ direction—which becomes proportional to $\alpha\,k^{2}_{KM}-\beta\,k^{4}_{\Gamma M}$, where $k_{ab}$ denotes the momentum component along direction abab—the FS becomes significantly flattened. As a consequence, the perfect nesting at transfer momentum $M_{i}$, which is typical for ordinary VHS in hexagonal lattice systems, is lifted. Although the majority of the DOS remains concentrated at the $M$-points, the band flattening induces an enhanced spread along the Fermi surface (FS) surrounding $M$.
Figure 1: Breathing kagome model. (a) Breathing kagome lattice configuration with hoppings up to fourth nearest neighbor. (b) Corresponding band structure with density of states, characterized by an algebraic divergence at the Fermi level. The inset displays the broadened Fermi surface at T=0.01tT=0.01\,t.
@@
Figure 2: Leading eigenvalues and order parameter. (a) Leading eigenvalues of the effective scattering vertex in the particle-hole channel on a transfer momentum $\boldsymbol{q}$-mesh in the Brillouin zone. The dominant eigenvalues are situated at the two inequivalent $K$-points. (b) Order parameter obtained by combining functional renormalization group calculations with Landau-Ginzburg and mean-field analysis.
While the pristine kagome Hubbard model, featuring perfect nesting and sublattice interference at its Van Hove fillings, has been studied extensively using FRG methods [55, 117, 91], the influence of HOVHS has mainly been investigated either by mean-field theories [136], or by renormalization group approaches applied to simple models known as patch theories [21, 34, 39, 123]. Relying on the assumption that the relevant fermions are those where the density of states is largest, these simple models discard the Fermi surface away from the vicinity of the HOVHS. For HOVHS near $M$, the FS scattering processes are then restricted to transfer momenta $\boldsymbol{q}=(M,\,\Gamma)$. However, by relaxing the constraint to transfer momenta which relate HOVHS points within our FRG analysis, we observe an unprecedented ordering tendency towards $\bm{q}=K$ density waves—signified by the dominant peaks at $K$ in Fig. 2a—with significant bond character, i.e. Kekulé order. Even though this wavevector is unrelated to the position of the HOVHS, we observe the Kekulé instability in an extensive parameter regime (see SM [106]).
This instability of the particle-hole channel breaks translation symmetry, and corresponds to the two inequivalent wavevectors $K$ and $K^{\prime}$ related by time-reversal symmetry (Fig. 2a). The enlarged $\sqrt{3}\times\sqrt{3}$ periodicity triples the size of the unit cell (Fig. 2b, black). The resulting order parameter
$\Delta^{\bm{q}}_{o_{1}o_{2}}(\bm{k})=\frac{1}{N_{\bm{k}}}\!\sum_{\bm{k}^{\prime}o_{3}o_{4}}\!V^{D}_{o_{1}o_{2}o_{3}o_{4}}(\bm{q},\bm{k},\bm{k}^{\prime})\langle{\hat{c}^{\dagger}}_{\bm{k}^{\prime}-\bm{q}o_{3}}{\hat{c}^{\vphantom{{\dagger}}}}_{\bm{k}^{\prime}o_{4}}\rangle,$ (2)
is given by the eigenstate of the two-particle vertex $V^{D}$, written in the direct particle-hole channel ($D$ channel), and corresponds to the $E^{\prime}_{3}$ irreducible representation of the extended point group $C^{\prime\prime}_{3v}$ (see SM [106]).
@@
$\Delta_{\boldsymbol{k}}(A,\phi)=A\,\big[e^{i\phi}\Delta_{\boldsymbol{k}}^{K}+e^{-i\phi}\Delta_{\boldsymbol{k}}^{{K^{\prime}}}\big]\,$, (3)
@@
It is noteworthy that our kagome Kekulé phase exhibits real-space features similar to those previously reported in honeycomb models. The VHS near the Fermi level in Co3Sn2S2 is the remnant of the so-called mixed ($m$)-type VHS in the undistorted lattice [74], near which the wavefunction at each $M$-point resides on two of the three kagome sublattices. The local density of states for such a Fermi surface in fact forms a honeycomb lattice, situated at the centers of the corner-sharing triangles [44]. Upon triggering a breathing distortion, the density of states depletes on one sublattice of this emergent honeycomb lattice; the result is that the Fermi level density of states near the HOVHS forms a partially sublattice-imbalanced honeycomb lattice, as can be seen in the nematic density of states simulated in Ref. [74]. In real-space, our Kekulé state therefore closely resembles the Kekulé orders found in honeycomb systems, but with an imbalance between the $A$ and $B$ sublattices. This correspondence can also be visualized through the line graph reconstruction of the kagome lattice, in which the lattice sites can be associated with the bonds of the emergent honeycomb lattice. This construction reveals a Kekulé-Y order and implies that the state given in Fig. 2b transforms in the same irreducible representation (see SM [106]).
Diffuse nesting.—The emergence of $K$-order across a broad interaction parameter regime indicates that the initial structure of the interaction vertex is largely irrelevant to the underlying ordering mechanism, implying that its origin can be discerned from a kinetic viewpoint, i.e. the identification of a generalized nesting condition. We consider the static particle-hole susceptibility at finite momentum and temperature [23],
@@
where $G_{o_{1}o_{2}}(\bm{k},\omega_{n})$ is the single particle propagator with momentum $\bm{k}$ and fermionic Matsubara frequency $\omega_{n}$ and $o_{i}$ indicates the sublattice degree of freedom. Figure 3a shows that at low temperature, the susceptibility peak at $\bm{q}=\Gamma$ dominates—in line with the expectation that translationally invariant orders should dominate at weak coupling. Yet at finite FS broadening, which in the kinetic model can be modeled with temperature (see Section SM 4 [106] for a more detailed discussion of FS broadening), the susceptibility at the transfer momentum $\bm{q}=K$ exceeds that at $\bm{q}=\Gamma$.
@@
One can understand Fermi surface nesting as the overlap of the FS with itself upon shifting by a momentum $\bm{q}$. For an unbroadened FS (zero temperature, no interactions/disorder), no significant overlap can be discerned for nonzero $\boldsymbol{q}$ nesting configurations, owing to the different curvatures of overlapping portions of the FS. This changes for a broadened FS, which we realize by introducing a finite temperature $T$ that leads to “smeared out” states near the Fermi level (Fig. 1b, inset) due to Fermi-Dirac statistics. For an exemplary broadening of $T=10^{-2}\,t$, Fig. 3b shows nested circular regions (black) that emerge when the broadened FS is shifted by $\boldsymbol{q}=K$, revealing the origin of the dominant finite-momentum peak in the susceptibility displayed in Fig. 3a. A detailed discussion and visualization for various levels of broadening and nesting vectors is provided in the SM [106].
@@
We emphasize that this argument generalizes beyond our finite temperature calculations, and is expected in the presence of other kinds of Fermi surface broadening, such as that due to interactions or disorder. Finite interactions of intermediately correlated systems allow FS scattering processes to access an increasing energy window around the Fermi level comparable to the broadening in Fig. 3b. As a consequence, we observe $K$-order across the predominant part of the finite interaction parameter regime within our FRG analysis (see SM [106]).
@@
Previous studies suggest enhanced nematic ordering tendencies associated with HOVHS [21, 34, 74]. These prior models are valid in the strict weak coupling regime, and indeed we observe competing $\boldsymbol{q}=\Gamma$ Pomeranchuck fluctuations—reflected by the peak at $\Gamma$ in Fig. 2a—which become dominant for very weak interaction parameters (FRG results shown in Section SM 3 [106]). Note that the enhanced DOS of HOVHS effectively places the system in an intermediate-coupling regime even for relatively small interactions—further disfavoring the applicability of previous approaches. This renders the functional renormalization group—valid from the weakly to intermediately coupled limits—a vital tool to appropriately capture the leading ordering tendency.
Figure 3: Susceptibility and Fermi surface nesting in the kinetic breathing kagome model with FS broadening modeled by temperature. (a) Maximal eigenvalue of the particle-hole susceptibility Eq. 4 for three different temperatures $T$. At higher $T$, the peak at momentum $K$ surpasses the $\Gamma$ peak. (b) Overlap (black) of initial (gray) and $K$-shifted (purple) Fermi surface for $T=10^{-2}\,t$. Circular nesting features drive $K$-order at finite $T$.
@@
correspondingly decreases, allowing for more localized Wannier functions [134, 121] (see Section SM 8 [106] for details and a quantitative illustration). This allows for a conceptually generalized model consisting of a piecewise series-expanded dispersion around the $M$-points of a hexagonal BZ with quartic dispersion along the $\Gamma M$ direction, as visualized in Fig. 4a. The shaded triangular regions represent the sections of the BZ where the individual expansions govern the dispersion. This piecewise definition preserves the lattice symmetry and ensures a continuous spectrum [106]. The resulting bandstructure of the effective single orbital model and its FS are displayed in Fig. 4b. Apart from the cusps at the $\Gamma K$ line, the FS retains the same directional band flattening features around the $M$-points. Consequently, diffuse nesting persists, and the dominance of $K$-order is observed within our unbiased FRG analysis, generalizing our proposed diffuse nesting mechanism to arbitrary lattice systems featuring such HOVHS.
@@
SM 1 Kinetic models with higher-order Van Hove singularities